OpenStudy (readergirl12):

Which of the ordered pairs in the form (x, y) is a solution of this equation? 5x-y/3=13 (2,-9),(3,-6) A. The first is a solution, but the second is not. B. The first is not a solution, but the second is. C. Both are solutions. D. Neither is a solution.

OpenStudy (calculusxy):

\[5x - \frac{ y }{ 3 } = 13 \rightarrow - \frac{ y }{ 3 } = 13 - 5x \rightarrow (3)(-y) = 3(13 - 5x) \rightarrow -3y = 39 -15x\] \[-3y = 39 -15x \rightarrow -y = -13+ 5x \rightarrow y = 13 - 5x \]

OpenStudy (calculusxy):

@AlexandervonHumboldt2 Can you just check whether this is correct?

OpenStudy (alexandervonhumboldt2):

you just need to substitute each point into the equation and solve .

OpenStudy (calculusxy):

so \(y = -5x + 13\) is correct?

OpenStudy (calculusxy):

don't you multiply both sides by 3 to get rid of it from y? the inverse of division is multiplication?

OpenStudy (calculusxy):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

woops yeah sorry

OpenStudy (alexandervonhumboldt2):

5x -y/3=13 -y/3=13-5x -y=(-5x+13)*3 y=15x-39

OpenStudy (calculusxy):

why did you not multiply 3 with -y?

OpenStudy (calculusxy):

@AlexandervonHumboldt2